# A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data

### v1

@@ Appendix B Synthetic Dimension Datasets: Details
Table 3: Descriptives for Synthetic Dimension Datasets: Sentence Counts, Sentence Lengths, and Total Generation Cost.
@@
Note: M = Mean Sentence Length of Dataset. Neutral = Neutral, unaltered, input sentences. Increase = Increase on the Dimension of interest. Decrease = Decrease on the Dimension of interest.
Figure 5: Distribution of Sentences in Each Sample of 50 Sentences.
@@ Appendix C In-Context Learning Paradigm
The study generated synthetic datasets to simulate changes in Sentiment and Intensity using 36,151 and 39,896 neutral baseline sentences, respectively. Neutral sentences were sampled by linking words in each sentence with their mean valence or arousal scores from the NRC-VAD lexicon (0-1) Mohammad (2018) and filtering by a dynamic range. This neutral range is adjusted from the median of each dataset by ±0.01, targeting 25th-75th percentile bounds or 500-1500 unique sentences per epoch. See Figures 7 and 7 for a breakdown of neutral sentence counts per epoch provided as input to the LLM using the prompts below.
@@
The ChatGPT API with a temperature setting of 1.00 was used to ensure semantic accuracy and prevent errors (Periti and Montanelli, 2024b), while allowing for a balance between deterministic and creative responses. Note that there were challenges in maintaining target terms in the sentences, particularly for positive sentiment variations. Fewer manual adjustments were needed for Intensity than Sentiment. GPT-4o struggled to vary 97% of the sentences to contain more positive sentiment for abuse (28), anxiety (110), depression (46), mental_health (1), trauma (2) as it replaced targets with positive terminology against instructions. For Intensity data, fewer sentences required manual alteration: only for abuse (4), depression (2), mental_health (2), trauma (1). Rows (196) were detected and manually altered to retain the target term while ensuring variation in the dimension relative to the neutral sentence. The final validated datasets, detailed in Table 6, are available on the GitHub repository: [MASKED LINK].
@@
mental_health Sentiment 6,348 6,348 6,348 21
@@
mental_illness Sentiment 2,552 2,552 2,552 9
@@ Appendix F List of Donor Terms: Synthetic Breadth
[h] [Uncaptioned image] Counts of synthetic sentences (donor-sibling contexts). Follow this GitHub link to access the counts of synthetic sentences for each five-year interval and the ranked lists for each sampling strategy (Boostrapped and Five-Year): [MASKED LINK]
Appendix G Multilevel Modeling Approach
To analyze the predictive effects of synthetic injections while accounting for hierarchical dependencies, we employ multilevel modeling (Gelman and Hill, 2007). These mixed linear models are well-suited for analyzing nested data, as they allow for the inclusion of fixed effects (e.g., injection_level) and random effects (e.g., variability across target). This approach leverages the full dataset while accounting for group-level structure, avoiding overfitting and unreliable estimates often encountered in simple linear regression when data points per group are limited.
Model Specifications
Null Model: To assess the necessity of incorporating random effects into the analysis, we initially fit a null model. This null model includes only a fixed intercept ($\beta_{0}$) and random intercepts ($u_{j}$) to account for variability across groups (target), and is represented as:
$y_{ij}=\beta_{0}+u_{j}+\epsilon_{ij},$
where $y_{ij}$ is the outcome variable (e.g., avg_valence_index_positive) for observation $i$ within group $j$, $u_{j}\sim N(0,\sigma_{u}^{2})$ capturing group-level random intercepts, where $\epsilon_{ij}\sim N(0,\sigma_{\epsilon}^{2})$ represents residual variability. The intraclass correlation coefficient (ICC) is calculated to quantify the proportion of variance explained by the grouping. ICC values exceeding 0.05 indicate meaningful variability, thereby justifying the inclusion of random effects. For this dataset, the ICC is calculated as:
$\text{ICC}=\frac{\sigma_{u}^{2}}{\sigma_{u}^{2}+\sigma_{\epsilon}^{2}},$
where $\sigma_{u}^{2}$ is the variance of the random intercepts and $\sigma_{\epsilon}^{2}$ is the residual variance. Full Model: Next, we fit a full model, which incorporates the fixed effect of injection_level ($\beta_{1}$) alongside the random intercepts, expressed as:
$y_{ij}=\beta_{0}+\beta_{1}\cdot\texttt{injection\_level}_{ij}+u_{j}+\epsilon_{ij}.$
This full model allows us to evaluate the predictive influence of injection_level while accounting for hierarchical dependencies in the data. Random Slopes Model: To further explore whether the effect of injection_level varied significantly across target, we tested an additional model with random slopes ($u_{j1}$) for injection_level, expressed as:
$\textstyle y_{ij}$ $=\beta_{0}+\beta_{1}\cdot\texttt{injection\_level}_{ij}+u_{j}$
$\quad+u_{j1}\cdot\texttt{injection\_level}_{ij}+\epsilon_{ij}.$
Here, $u_{j1}\sim N(0,\sigma_{u1}^{2})$ represents the variability in slopes across groups.

### v2

@@ 3.2 Evaluation Framework
Table 2: Descriptive Statistics for Synthetic Dimension Datasets: Sentence Counts, Mean Lengths, and Total Generation Cost.
@@
Note. M = Mean sentence word length. Neutral = Original input sentences. Increase = Sentences modified to increase the dimension of interest. Decrease = Sentences modified to decrease the dimension of interest.
Experimental Settings:
We implemented two sampling strategies to assess the sensitivity of LSC detection methods. Bootstrap sampling involved randomly selecting 50 sentences with replacement from the full corpus (natural or synthetic), repeated over 100 iterations. This allowed any sentence to be sampled multiple times, enabling robust estimation through resampling. Five-year interval sampling selected up to 50 unique sentences per time bin, repeated 10 times. Sentences were not repeated within an iteration but could appear across iterations. This approach ensured balanced temporal coverage and better reflects natural language variability over time. Each sampling strategy had a control condition, where sentences were shuffled to balance natural and synthetic ones per sample, providing a baseline to isolate the effect of synthetic injection, following prior work in computational linguistics (Dubossarsky et al., 2017, 2019).
For each sampling strategy, synthetic sentences were injected into (up to) 50-sentence samples at increasing proportions: 20%, 40%, 60%, 80%, and 100% (see Figure 2). These injection levels simulate increasing semantic saturation to test method sensitivity (Stage 2). For method selection (Stage 3), only the 0% and 100% injection levels were used. This enables the clearest contrast between unchanged (fully natural) and fully altered (synthetic) samples, maximizing the signal-to-noise ratio when evaluating which method detects the greatest magnitude of semantic change.
Figure 2: Distribution of Sentences in Each Sample.
3.3 Synthetic Dataset Generation
For each semantic dimension and target (Section 3.1.3), we generated datasets by inducing change in natural sentences drawn from the corpus (Section 3.1.1). Each dataset included up to 1,500 randomly sampled sentences per five-year interval used in the experimental settings (see Section 3.2). See Appendix B for examples of synthetic sentences for each dimension and target.
3.3.1 Sentiment and Intensity
To generate the synthetic Sentiment and Intensity datasets, we employed few-shot ICL with GPT-4o to vary these dimensions. First, neutral diachronic sentences from the corpus (see Section 3.1.1) were sampled as outlined in Appendix C. Second, a psychology scholar crafted five Chen et al. (2023) diverse sentence variations for each target, to serve as few-shot demonstrations for the LLM in the prompt outlined below. This prompt includes construct definitions, a task description, and few-shot examples to generate theory-driven change — that is, changes aligned with the dimensions in the SIBling framework (e.g., shifts from positive to negative sentiment or from high to low intensity). For ‘scholar-in-the-loop’ few-shot demonstrations, see Appendix D (Sentiment), Appendix E (Intensity), and the few-shot demonstration examples below.
Third, the prompt was refined during pilot tests (varying 10 input sentences for each target). Fourth, for each of the neutral sentences (Sentiment: 36,151; Intensity: 39,896), we made one inference call to GPT-4o through the OpenAI API to generate variations of Sentiment (positive/negative) or Intensity (high/low). Fifth, validation checks required checking if output sentences retained the target term. A few manual adjustments were made by a tertiary-educated English native speaker to ensure the final sentences retained the target in a similar location to the original input sentence. Few output sentences required manual adjustment in the final datasets: 0.25% for the synthetic Sentiment dataset, and 0.01% for the synthetic Intensity dataset. See Table 2 for the counts of input/output sentences and final prompts (dataset costs: for Sentiment = 115 \$US; for Intensity = 136 \$US).
3.3.2 Breadth
Unlike Sentiment and Intensity, current Breadth measures have no score that assigns a mid-point with which to obtain neutral sentences to vary. Therefore, to induce semantic breadth, we adapted Dubossarsky et al.’s (2019) replacement strategy, using WordNet 3.0 (see Section 3.1.2) to expand a target word’s usage by incorporating contexts from donor terms, thereby broadening its semantic range without altering its core meaning. Relevant synsets were identified and filtered for psychological relevance using keyword matching [^5] Psychology key terms: ”abnormality”, ”abnormally”, ”emotional”, ”feeling”, ”feelings”, ”harm”, ”hurt”,”mental”, ”mind”, ”psychological”, ”psychology”, ”psychiatry”, ”syndrome”, ”therapy”, ”treatment”. and semantic similarity thresholds. Donor terms (co-hyponyms with the target) were filtered using Lin similarity (0.5) [^6] Information content values from the psychology corpus. and cosine similarity (0.7) using embeddings from BioBERT Lee et al. (2020), a pre-trained language model for biomedical text mining, to capture context-dependent meanings of synset glosses in 768-dimensional vectors. See Appendix F for the full list of siblings (i.e., co-hyponyms). The sibling replacement process identifies and replaces sibling terms with the target, as illustrated below. To sample representatively from the sibling list, a round-robin strategy was used, where the algorithm runs through each sibling in the list in batch sizes, randomly sampling up to 50 unique sentences before proceeding to the next one. In this way, it exhaustively samples up to 1,500 unique sentences per epoch and injection level to create the final synthetic breadth dataset.
Dataset Creation for Synthetic Breadth Replacement Strategy: Randomly sample sentences containing co-hyponyms of the target term from the validated list and replace the co-hyponym with the target to be used as a synthetic sentence. ————————————————————— [One example for mental_health below.] Donor Context: The ‘Angry and Impulsive Child’ and ’Abandoned and Abused Child’ modes uniquely predicted dissociation scores. Synthetic Context: The ‘Angry and Impulsive Child’ and ’Abandoned and Abused Child’ modes uniquely predicted mental health scores.
3.4 Quantifying Lexical Semantic Change
3.4.1 Semantic Dimensions
We applied methodologies from Baes et al. (2024) to assess changes in semantic Sentiment, Intensity, and Breadth. Sentiment and Intensity were quantified by assigning ordinal valence and arousal scores, based on Warriner et al. (2013) norms, matched to collocates within ±5 words of the target, using a scale (ranging from 1-9) from ‘extremely unhappy’ to ‘extremely happy’ for valence, and ‘extremely calm’ to ‘extremely agitated’ for arousal. Scores were frequency-weighted, normalized, and averaged across bins, yielding indices from 0 (negative/low arousal) to 1 (positive/high arousal). Breadth was measured by averaging the cosine distances between sentence-level embeddings from the SentenceTransformer model ‘all-mpnet-base-v2’ [^7] Microsoft pretrained network (109M model params) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (MPNet). The Breadth score indicates semantic range variation from 0 (no change) to 1 (maximum variation). See Appendix G.1.1 for further details.
For comparative analysis, Sentiment was compared against the Deberta-v3-base-absa-v1.1888yangheng/deberta-v3-base-absa-v1.1 (184M model params): https://huggingface.co/yangheng/deberta-v3-base-absa-v1.1 aspect-based sentiment analysis (ABSA) classification model, which we adapted to output continuous sentiment scores ranging from 0 (fully negative) to 1 (fully positive). Because Intensity is a novel dimension, it lacks comparable methods. For the Breadth comparisons, we used Cassotti et al.’s (2023) SOTA Periti and Tahmasebi (2024) word transformer "XL-LEXEME" [^9] XL-LEXEME (~550M model params) https://huggingface.co/pierluigic/xl-lexeme (XLL). While MPNet generates sentence embeddings through pooling tokens, which dilutes word-specific information, XLL uses a bi-encoder architecture that focuses on word-specific attention, using polysemy as a proxy for meaning divergence during training. See Appendix G.1.2 for further details on the quantification of semantic Breadth.
3.4.2 General Lexical Semantic Change
To quantify general LSC, we used the SOTA LSC score (Cassotti et al., 2023), which calculates the Average Pairwise Cosine Distance (APD) between sentence embeddings from two time periods, following the approach of Giulianelli et al. (2020). We extended this method to compare embeddings from different bins within the same iteration. See Appendix G.2 for more details.
3.5 Statistical Strategy
To test whether synthetic change influences SIB scores (RQ1), we fit separate mixed linear models for each outcome variable (Valence, Arousal, and Breadth), with random intercepts for target terms and fixed effects for injection level (see Appendix H for model specifications). To assess the sensitivity of various computational methods to detecting LSC (RQ2), we calculated the percent relative change in a target word’s context between the 0% (fully natural) and 100% (fully synthetic) injection levels, as defined in Equation 1:
$\Delta\%=\frac{X_{100}-X_{0}}{X_{0}}\times 100$ (1)
Here, $X_{0}$ and $X_{100}$ denote the score at 0% and 100% injection, respectively. This index quantifies how much the contextual meaning of the target word shifts when all natural sentences are replaced with synthetic ones.
For the XLL LSC Score, we further normalized this change to account for internal variability within each bin, as shown in Equation 2. We used the APD, the mean cosine distance between all sentence pairs, [^10] APD reflects semantic dispersion; higher APD implies greater variability. to control for within-bin noise:
$\Delta=\frac{\text{APD}(X_{100}\text{-between-}X_{0})}{\max[\text{APD}(X_{0}\text{-within-}X_{0}),\text{APD}(X_{100}\text{-within-}X_{100})]}$ (2)
The numerator reflects divergence between bins, while the denominator accounts for within-bin variability, ensuring that the score reflects genuine LSC signal rather than internal fluctuation.
4 Results
Synthetic Change Effects:
To address RQ1, we tested whether SIB scores from Baes et al.’s (2024) SIBling framework are associated with levels of synthetically induced change. The hypothesis was supported: standardized SIB scores across all three dimensions — Sentiment, Intensity, and Breadth — demonstrated consistent increases or decreases as a function of injection level, across targets and both sampling strategies: bootstrap (Figure 3) and five-year intervals (Figure 4).
Figure 4 illustrates SIB scores by five-year intervals across injection levels and conditions for three of the six targets. The black line in each panel depicts the changes in SIB scores over time in the natural corpus, while the colored lines represent injection levels. The intensity of the color reflects the percentage of synthetic sentences injected into each sample (20% to 100%). As intended, injection levels altered the height, but not necessarily the slope, of SIB scores on the y-axis, indicating changes in the magnitude of the score rather than trend over time. For breadth (bottom panel), only upward (broadening) changes were modeled.
Mixed-effects linear models confirmed that a one-standard-deviation increase in injection level significantly predicted changes in SIB scores in the intended direction (Table 3; see also Appendix H).
Score Valence Arousal Breadth
$\beta^{+}$61* .64* .43*
$\beta^{-}$-.31* --.64*
Table 3: Coefficients of random intercept (by target) models predicting SIB scores from injection level.
Control Experiments:
As illustrated in Figure 3, controlling for synthetic injection level by re-analyzing data with shuffled sentences for uniform distribution revealed flat SIB score trends in bootstrapped settings. Appendix I illustrated how, in the control condition for the five-year interval setting, SIB scores in those samples tend to converge to a midpoint between natural and synthetic data.
Comparative Method Evaluation:
To address RQ2 — identifying which methods are most sensitive to synthetically induced changes in SIB — we compared the performance of alternative change detection methods across each dimension. Results were mixed. For Sentiment, both the Valence index and ABSA’s Sentiment score were sensitive to detecting synthetic change, though ABSA outperformed the Valence index in 10/12 cases. For Intensity, the Arousal index shows sensitivity to detecting variations in synthetic Intensity. For Breadth, XLL outperformed MPNet in 4/6 cases using the Breadth score.
Critically, XLL-LSC score is completely insensitive to detecting changes in either Sentiment or Intensity. XLL-LSC can only indicate change via positive change values, while negative values indicate that the within-bin variance is greater than the change scores between bins. See Appendix J for between- and within-bin LSC scores across all synthetic injection levels. Thus, the negative scores observed in Sentiment and Intensity (except for mental illness) establish that XLL was unable to detect any change signal in these words. XLL-LSC detects changes in synthetic Breadth for 2/6 terms.
5 Discussion
The present study introduced LSC-Eval, a three-stage general evaluation framework that: (1) creates synthetic datasets featuring scholar-in-the-loop LLM-generated sentences to simulate various kinds of LSC; (2) leverages these datasets to assess the sensitivity of computational approaches to synthetic change; and (3) evaluates the suitability of these methods for specific dimensions and domains. We applied this framework to generate synthetic datasets that induce changes across the three dimensions (Sentiment, Intensity, and Breadth) of a novel multidimensional LSC framework, SIBling Baes et al. (2024), using examples drawn from psychology to compare the suitability of alternative methods for detecting these synthetic changes.
Importantly, these synthetic datasets reflect theory-driven change, operationalized through construct definitions and distinctions drawn from the SIBling framework (Baes et al., 2024), which integrates established traditions in lexical semantics, connotation, and affect. This principled approach enables interpretable and targeted benchmarking, and lays the groundwork for generalizing LSC-Eval to other theoretical models and domains.
Our findings support the hypothesis that recently proposed methods (Valence index, Arousal index, Breadth score; Baes et al., 2024) detect synthetic changes on the SIB (i.e., Sentiment, Intensity, and Breadth) dimensions. This directly addresses RQ1, which asked whether synthetic datasets can be used to validate methods for measuring LSC dimensions. Control analyses, which adhered to computational linguistics standards Dubossarsky et al. (2017, 2019), confirmed the absence of these effects in shuffled controls. The implications of these findings are two-fold. First, the ability of SIB methods to detect changes introduced via silver-standard synthetic data validates their sensitivity and reliability for detecting and quantifying variation in SIB, even in artificial environments. Second, it confirms the validity of using LLM-generated sentences in our ICL evaluation suites.
To address RQ2 — identifying which computational methods are most sensitive to synthetically induced changes in SIB — we demonstrated how a synthetic change detection task can assess the sensitivity of various approaches, guiding the selection of the most suitable method for specific dimensions and domains. Baes et al.’s (2024) tools, which validated the synthetic SIB datasets, were supported by alternative methods that consistently detected synthetic changes across conditions and targets, providing further validation. The Valence index and ABSA’s Sentiment score detected variations in synthetic Sentiment, while Breadth scores derived from XLL and MPNet models detected increases in synthetic Breadth. These results suggest that the NLP-based ABSA method is generally more sensitive to synthetically induced change than Warriner-based approaches, which rely on static Valence and Arousal ratings. Future empirical studies on Sentiment and Breadth may consider adopting these NLP-based models, either as replacements for, or in addition to, traditional lexicon-based methods.
Notably, when computing the general LSC score using the SOTA LSC model XLL Cassotti et al. (2023), it was not sensitive to detecting Sentiment and Intensity. Although XLL shows some sensitivity to identifying synthetic increases in Breadth, it registers a more substantial change when the Breadth score is adjusted according to the method introduced by Baes et al. (2024). It uses the within-bin average cosine distance of target-containing sentences as a proxy for the expansion (broadening) or contraction (narrowing) of a word’s contextual usage. The inability of XLL to detect the affective dimensions of LSC highlights the necessity of evaluating SOTA models before deploying them in new domains. Future research should investigate whether this weakness in detecting affective dimensions is specific to XLL or extends to other contextualized models in other corpora and concepts. This inquiry is particularly salient given recent advances in analyzing fine-grained, continuous semantic shifts through “diachronic word similarity matrices using fast and lightweight word embeddings over arbitrary time periods" Kiyama et al. (2025). Our results support the idea that current SOTA methods are only state-of-the-art with respect to a particular kind of change or a specific domain and should not be assumed to generalize beyond them.
6 Conclusion
The current study introduced a novel general-purpose evaluation framework, LSC-Eval, for evaluating methods that detect LSC. Its three-stage approach: 1) generates LLM-based synthetic datasets with silver-standard labels that simulate changes in kinds of LSC; 2) uses these datasets to evaluate the relative sensitivity of computational methods in a synthetic change detection task; and 3) identifies the most suitable method for detecting synthetically induced changes across specific dimensions and domains. We applied this framework to a set of psychological terms. Findings not only supported the validity of proposed computational methods for measuring changes in SIB, but also established a controlled experimental standard for rigorously evaluating existing LSC detection methods and exploring alternative computational approaches. This work is crucial for addressing the substantial gap created by the lack of historical benchmark datasets, which has previously hindered the standardization of metrics and fair comparison of methods. While this framework benefits all disciplines (e.g., biomedicine, law, theology), LSC-Eval is particularly valuable in the social sciences and humanities, where domain-specific constructs require specialized tools for capturing nuanced semantic change.
Limitations
Limitations inform future directions. Evaluating the quality of LLM prompt and demonstration examples in the few-shot ICL paradigm is challenging. As LLM evaluation standards are developed Chang et al. (2024); Ziems et al. (2024), future research might explore automated strategies such as updating prompts based on examples (DSPy) [^11] https://dspy.ai or comparing LLM output from different prompts using a free, unified interface. [^12] https://github.com/marketplace/models/azure-openai/gpt-4o/playground LLM choice in the evaluation pipeline could be expanded to include open-source models (e.g., FlanT5-XL, Mistral-7B, Mixtral-8x7B). Future work could also compare this approach to a non-generative rule-based one.
Furthermore, our study benefited from using GPT-4o, which is trained on US English and is therefore well-suited for analyzing texts within the Western-centric domain of psychology Varnum et al. (2024). However, cultural and linguistic biases of LLMs can pose challenges in adapting our evaluation pipeline to other languages Havaldar et al. (2023), although few-shot ICL has proven effective in low-resource languages Cahyawijaya et al. (2024). Despite the tendency of LLM training data to skew towards the recent past, manual inspection of results demonstrate the successful generation of quality sentences that spanned a 1970 to 2019 time period. Future work should focus on refining these models to broadly apply across cultural contexts, languages, and historical periods.
Furthermore, the present study does not specify which sense of the term is semantically expanded. Attempting to integrate senses into the synthetic data generation pipeline may provide richer insights. While the specialized psychology corpus and target words exhibit limited senses, general domain corpora introduce ambiguous contexts (e.g., the economic sense of “depression"). Further research is needed on sense-specific LSC.
Ethical Considerations
Acknowledgments
We acknowledge the support of the University of Melbourne’s general-purpose High Performance Computing system, Spartan Lafayette et al. (2016), which provided the computational resources necessary for efficient embedding encoding using transformer models. We also thank Jeremy Silver from the Statistical Consulting Centre for his valuable feedback on the random effects modelling.
This research was supported by an Australian Government Research Training Program Scholarship and funded, in part, by Australian Research Council Discovery Project DP210103984 and the research program “Change is Key!", supported by Riksbankens Jubileumsfond (under reference number M21-0021).
Appendix A Corpus Counts of Target Terms
Appendix B Examples of Synthetic Sentences for each Dimension and Target
@@ Appendix C In-Context Learning Paradigm
The study generated synthetic datasets to simulate changes in Sentiment and Intensity using 36,151 and 39,896 neutral baseline sentences, respectively. Neutral sentences were sampled by linking words in each sentence with their mean valence or arousal scores from the NRC-VAD lexicon (0-1) Mohammad (2018) and filtering by a dynamic range. This neutral range is adjusted from the median of each dataset by ±0.01, targeting 25th-75th percentile bounds or 500-1500 unique sentences per epoch. See Figures 8 and 8 for a breakdown of neutral sentence counts per epoch provided as input to the LLM using the prompts below.
@@
The ChatGPT API with a temperature setting of 1.00 was used to ensure semantic accuracy and prevent errors (Periti and Montanelli, 2024b), while allowing for a balance between deterministic and creative responses. Note that there were challenges in maintaining target terms in the sentences, particularly for positive sentiment variations. Fewer manual adjustments were needed for Intensity than Sentiment. GPT-4o struggled to vary 97% of the sentences to contain more positive sentiment for abuse (28), anxiety (110), depression (46), mental_health (1), trauma (2) as it replaced targets with positive terminology against instructions. For Intensity data, fewer sentences required manual alteration: only for abuse (4), depression (2), mental_health (2), trauma (1). Rows (196) were detected and manually altered to retain the target term while ensuring variation in the dimension relative to the neutral sentence. The final validated datasets, detailed in Table 6, are available at: https://github.com/naomibaes/Synthetic-LSC_pipeline.
@@
mental health Sentiment 6,348 6,348 6,348 21
@@
mental illness Sentiment 2,552 2,552 2,552 9
@@ Appendix F List of Donor Terms: Synthetic Breadth
[h] [Uncaptioned image] Counts of synthetic sentences (donor-sibling contexts). Follow this GitHub link to access the ranked lists for each sampling strategy (Boostrapped and Five-Year): https://github.com/naomibaes/LSCD_method_evaluation/tree/main/supplementary_materials/synthetic_breadth_siblings
Appendix G Quantifying Lexical Semantic Change
G.1 Semantic Dimensions
G.1.1 Sentiment and Intensity:
[…]
G.1.2 Breadth:
To estimate the semantic broadening (expansion) or narrowing (contraction) of a word’s meaning, we calculate the average cosine distance between sentence-level embeddings of a target term, as in Baes et al. (2024). The SentenceTransformer model ‘all-mpnet-base-v2’ [^17] Microsoft pretrained network (109M model params) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is used to generate these embeddings. The Breadth score, $B$, is derived by averaging the cosine distances, $\delta$, across all unique pairs of sentence embeddings within each iteration, and then averaging these scores across all iterations within each bin, as in Equation 4:
Here, $\delta(s_{i,k}^{t_{j}},s_{j,k}^{t_{j}})$ calculates the cosine distance between two sentence embeddings in the same iteration $k$ in bin $t_{j}$. $N_{k}$ is the number of sentence embeddings in iteration $k$; $I_{j}$ is the number of iterations in bin $t_{j}$. Higher scores indicate greater variation in the target’s semantic range. Scores range from 0 (no variation) to 1 (max variation). The sentence transformer "all-mpnet-base-v2" (MPNet) from Cassotti et al. (2023) is compared with the SOTA transformer "XL-LEXEME" [^18] XL-LEXEME (~550M parameters): https://huggingface.co/pierluigic/xl-lexeme (XLL). MPNet pools tokens to produce sentence embeddings, which dilutes word-level information, whereas XLL employs a bi-encoder focused on word-specific attention, [^19] Only the first target occurrence is attended to. using polysemy as a proxy for semantic divergence during training (WiC; Pilehvar and Camacho-Collados, 2019).
G.2 General Lexical Semantic Change:
General lexical semantic change is evaluated using the LSC score shown in equation 5.
Appendix H Linear Mixed-Effects Modeling for SIB Dimensions
To evaluate the effects of synthetic injection, we fit mixed-effects linear regression models with fixed effects for injection_level and random intercepts for target. This approach accounts for repeated observations per target and treats the grouping variable as sampled from a population of effects. The grouping variable target had six levels, a sufficient number for reliable variance estimation.
Model Structure and Rationale:
We first fit a null model with only random intercepts to estimate the intraclass correlation coefficient (ICC), which quantifies variance attributable to group-level clustering. ICC values above 0.05 suggest meaningful clustering and justify random intercepts. Next, we fit a random intercepts model including injection_level as a fixed effect, expressed in Equation 6.
$y_{ij}=\beta_{0}+\beta_{1}\cdot\texttt{injection\_level}_{ij}+u_{j}+\epsilon_{ij}$ (6)
where $y_{ij}$ is the outcome for observation $i$ in group $j$, $\beta_{0}$ is the fixed intercept, $\beta_{1}$ is the fixed slope for injection level, $u_{j}\sim\mathcal{N}(0,\sigma_{u}^{2})$ is the group-specific random intercept, and $\epsilon_{ij}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is the residual error. We also tested a random slopes model allowing injection_level slopes to vary by target, but with only six groups, these models often failed to converge or showed negligible slope variance.